OpenStudy (anonymous):

2x + 1 greater than or equal to 5 or -3x > -9 Solve the inequality above. @mikeyymonroe

OpenStudy (anonymous):

i feel like i know how to do this but i forgot...

OpenStudy (mikeyymonroe):

so is there a value for x or is it just there for looks?

OpenStudy (anonymous):

lol yes XD

OpenStudy (mikeyymonroe):

what is the value?

OpenStudy (studygurl14):

\[2x + 1 \ge 5 \] Subtract 1 from both sides: \[2x \ge 4\] Divide by 2 \[x \ge 2\] or \[-3x > - 9\] Divide by -3. Remember to flip the sign \[x < 3\] Do you have to graph?

OpenStudy (anonymous):

im figuring it out

OpenStudy (studygurl14):

Well, anyway, if you don't have to graph, the answer is... x is greater than or equal to 2 OR x is less than 3

OpenStudy (anonymous):

thank you :)

OpenStudy (studygurl14):

yw :)